OpenStudy (anonymous):

If the quadratic equation is used to find the solution set of 3x^2 + 4x - 2 = 0, what are the solutions?

OpenStudy (anonymous):

x = 1/3 (sqrt(10)-2) or x = 1/3 (-2-sqrt(10))